Derivative: g[sup](1)[/sup]= 3x[sup]2[/sup]-4x+1 and graph of g passes through point P(2,3). Find g(x).
1
Expert's answer
2011-01-31T12:51:27-0500
Let's integrate g(1):
int(g(1)) = x3- 2x2 + x + C;
As the function passes through the point (2,3), we can find the constant C:
23- 2*22 + 2 + C = 3
C = 3-2 = 1
Then
g(x) = x3- 2x2 + x + 1.